Minutes of the Trigger Operation Meeting (, 2007)
Attendees: Debbie Bard, Fang Fang, Rahmat Rahmat.
- Paged last night because of TSF damage. odf log file shows spurious interrupt messages between all the out of frame, out of synch and out of error messages so I powercycled the crate. All was well afterwards, but I don't know what caused it.
- Over the weekend I tested the 3 spare TPBs, two of which were recently repaired. The IR2 spare was good and has been left in the crate, of the two repaired boards one was good and one showed damage.
- Correct L3 cosmics JAS reference plots were installed with the help of Jim - the procedure is different compared to LiveFastMon plots and I'll update the reference guide accordingly.
- Fang should have a chance to test the GLT libraries this afternoon, when they turn the magnet on again.
- In the meantime, she has been doing memory tests in IR2 - it works at the teststand but in IR2 it doesn't work at all. The framework is correct but it never gives the right answer, plus after the ROM is rebooted at the end of the tests the link for the zpd crate is lost and the crate must be powercycled (odf log file).
Data Quality and Monitoring
- See Rahmat's hypernews about a problem with run 70018 lumi ratio. It is significantly lower than all other runs in all lumi plots, but all other quantities are OK for it. Could it be because of low statistics, or is there more investigation needed?
This page is created by Debbie Bard (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.